OpenStudy (anonymous):

52. Increasing life expectancy. A smoker is 15 times more likely to die of lung cancer than a nonsmoker. An ex-smoker who stopped smoking t years ago is w times more likely to die of lung cancer than a nonsmoker, where w = 15 – t. Graph the equation and use the graph to estimate how much more likely it is for Sandy to die of lung cancer than Polly, if Polly never smoked and Sandy quit yr ago.

OpenStudy (anonymous):

Can somebody please help me with this

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

did u try graphing 15-t

OpenStudy (anonymous):

first equation is smoker to non smoker y=15t second is quitter to non smoker w=15-t y=wx y=(15-t)x